OpenStudy (anonymous):

center at (-5,1) circle passes through (2,-1)

OpenStudy (gorv):

(x+5)^2+(y-1)^2=(-5-2)^2+(1+1)^2=49+2=51 (x+5)^2+(y-1)^2=51 r=distance between centre and point given